OpenStudy (anonymous):

A 4000kg lorry is moving with a constant velocity of 30ms^-1 . The driver of the lorry then decelerates with 0.5ms^-1 . Calculate the energy loss while the driver decelerates. Show working please.

OpenStudy (anonymous):

the deceleration for 10 s.

OpenStudy (anonymous):

m= 4000kg u= 30m/s v=0m/s a=-0.5m/s^2 \[v^2-u^2=2.a.s\] put the values and get 's'.. s=900m then, \[W= m.a.s\] 'W' is the amount of work done to stop the lorry=Energy loss now get the answer.. :)

OpenStudy (anonymous):

-tive value indicates energy loss..

OpenStudy (anonymous):

\[v=u+at\]\[v=30+(-0.5)(10)\]\[v=25\] energy lost=\[0.5mv^2 - 0.5mu^2\]=\[0.5(4000)(25^2 - 30^2) = -550000J\]

OpenStudy (anonymous):

@glgan1 You got the answer right . Why do u need to find v? i supposed the final velocity is 0? I dont understand this part? What formula is that? energy lost= 0.5mv2−0.5mu2

OpenStudy (anonymous):

sorry i had not noticed the tym given...

OpenStudy (anonymous):

nope. the final velocity isn't 0. it doesn't state that it stops. it just says that it decelerates for 10s so we can't assume it's stopped. 0.5mu^2 is the initial kinetic energy for the lorry and 0.5mv^2 is the final kinetic energy of the lorry which is after deceleration.

OpenStudy (anonymous):

@rahulchatterjee Its alright. :)

OpenStudy (anonymous):

@glgan1 Thank u

OpenStudy (anonymous):

you're welcome.